OpenStudy (studentchaitrali):

Pls pls pls pls help TRIGONOMETRY (CotA+cosecA-1)/(cotA-cosecA+1)=(1+cosA)/sinA Pls prove this...I can't do it.

OpenStudy (anonymous):

@dan815 @chmvijay

OpenStudy (triciaal):

review cos^2 + sin^2 = 1 cosec = 1/ sin cot = 1/tan 1 +tan^2 = sec^2

OpenStudy (triciaal):